OpenStudy (anonymous):

WORD PROBLEM - 2 equations & 2 variables: "A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). The one-time fixed costs will total $28,650. The variable costs will be $9.75 per book. The publisher will sell the finished product to bookstores at a price of $22.25 per book. How many books must the publisher produce and sell so that the production costs will equal the money from sales? Let B = # of books produced @Satellite73

hero (hero):

y = 28650 + 9.75x y = 22.25x y = y 28650 + 9.75x = 22.25x

hero (hero):

What's the matter? lost?

OpenStudy (anonymous):

I do not understand. Teacher said to use the variable B for # of books produced and set up 2 equations.

zepdrix (zepdrix):

Word problems? :O ughhh lol

OpenStudy (anonymous):

please help

hero (hero):

Well, swap variables x and y with A and B

OpenStudy (anonymous):

\(B\) is the number of books produced, then you will make \[22.25B\] dollars, at a cost of \[28,650+9.75B\] so figure out where they aer equal write \[2865+9.75B=22.25B\] and solve for \(B\)

OpenStudy (anonymous):

i honestly don't see how you get two variables out of this, unless you simply introduce another one, like \[C=2865+9.75B\] and \[P=22.25B\] then say set \(C=P\) but is is the same as what i wrote and what @hero wrote above